### v1

# TEXONO Collaboration

Abstract
After decades of experimental efforts, the DAMA/LIBRA($\rm DL$) annual modulation (AM) analysis on the $\chi{\rm N}$ (WIMP Dark Matter interactions on nucleus) channel remains the only one which can be interpreted as positive signatures. This has been refuted by numerous time-integrated (TI) and AM analysis. It has been shown that $\rm\chi e$ (WIMP interactions with electrons) alone is not compatible with the DL AM data. We expand the investigations by performing an AM analysis with the addition of $\rm\chi e$ long-range and short-range interactions to $\chi{\rm N}$, derived using the frozen-core approximation method. Two scenarios are considered, where the $\chi{\rm N}$ and $\rm\chi e$ processes are due to a single $\chi$ (${\Gamma}_{tot}^{1\chi}$) or two different $\chi$’s (${\Gamma}_{tot}^{2\chi}$). The combined fits with $\chi{\rm N}$ and $\rm\chi e$ provide stronger significance to the $\rm DL$ AM data which are compatible with the presence of additional physical effects beyond $\chi{\rm N}$ alone. This is the first analysis which explores how $\rm\chi e$ AM can play a role in DL AM. The revised allowed regions as well as the exclusion contours from the other null AM experiments are presented. All DL AM allowed parameter spaces in $\chi{\rm N}$ and $\rm\chi e$ channels under both ${\Gamma}_{tot}^{1\chi}$ and ${\Gamma}_{tot}^{2\chi}$ are excluded at the 90% confidence level by the combined null AM results. It can be projected that DL-allowed parameter spaces from generic models with interactions induced by two-WIMPs are ruled out.
There are compelling experimental evidence that about one-quarter of the energy density of the Universe is composed of dark matter (DM), whose exact nature and properties remain unknown. Searches of DM in numerous directions with diverse techniques are intense areas of fundamental research [1]. A favored DM candidate is the weakly interacting massive particle (WIMP, denoted as $\chi$[2, 3]. Direct experimental searches assume finite interactions between WIMPs with electrons ($\rm\chi e$) and nuclei ($\chi{\rm N}$). Positive signatures manifest as excess events over known background in the measured “time-integrated” (TI) energy spectra and in their annual modulation (AM) due to changes of the relative velocity between Earth and the WIMPs in the galactic halo [4]. TI analysis is sensitive to uncertainties of background modeling while AM analysis only requires the background is stable with time but independent of other details.
After decades of experimental efforts, the only result consistent with positive WIMP signatures is from the AM analysis on $\chi{\rm N}$ from the DAMA/LIBRA($\rm DL$) experiment with NaI(Tl) scintillating crystals [5]. This interpretation, however, is challenged and rejected by numerous $\chi{\rm N}$ experiments with TI analysis [1] and several with AM analysis [6, 7, 8, 9, 10] using a variety of targets. There are attempts to explain the $\rm DL$ AM data with scenarios other than $\chi{\rm N}$ detection, such as complications in the analysis procedures [11, 12]. In addition, it has been shown [13] -- and verified in our studies -- that AM analysis with the $\rm\chi e$-channels only (setting $\sigma^{\rm SI}_{\chi{\rm N}}{=}0$ in Eq. Dark Matter Annual Modulation Analysis with Combined Nuclear and Electron Recoil Channels) is unable to provide an acceptable fit to the $\rm DL$ data. The maximum electron recoil energy even for infinite $\rm m_{\chi}$ is about $3~{}\rm keV_{ee}$, in serious discrepancy with the data.
Figure 1: (a) Published AM amplitudes data from DM experiments with NaI in $\rm DL$ [14], COSINE [9] and ANAIS [15, 10], Xe in XMASS [6] and Ge in CDEX [7]. For display purposes, the depicted event rates are re-scaled to account for the differences in bin-sizes and target mass number squared. The expected differential AM spectra for positive WIMP signatures in: (b) $\rm\chi N^{SI}$ and (c) in $\rm\chi e^{LR}$ and $\rm\chi e^{SR}$ for all three targets taking the case of $\rm m_{\chi}{=}80~{}{\rm GeV}$ as example, together with the interference spectrum between $\rm\chi e^{LR}$ and $\rm\chi e^{SR}$ for NaI. The $\rm\chi e^{LR}$ and $\rm\chi e^{SR}$ spectra in (c) are superimposed as dotted lines in (b) f showing their responses to AM are very different from $\rm\chi N^{SI}$. The interference spectrum between $\rm\chi e^{LR}$ and $\rm\chi e^{SR}$ for NaI is also displayed in (c).
Figure 2: Photo-ionization cross-sections derived by FCA following Ref. [16] for (a) Na and (b) I, both as atoms and ions, within 1 $\rm keV_{ee}$ to 30 $\rm keV_{ee}$ energy transfer, The experimental data [17] and theoretical values [18, 19] are superimposed. Relative deviations from benchmark data are displayed.
We performed an analysis including both $\rm\chi e$ and $\chi{\rm N}$ interactions. The combined analysis would provide better matching statistically to the $\rm DL$ data. Three types of WIMPs interactions are included: spin-independent WIMP-nucleon interaction ($\rm\chi N^{SI}$), long and short-range WIMP-electron interaction ($\rm\chi e^{LR}$ and $\rm\chi e^{SR}$, respectively) [16]. Their cross-sections (respectively $\sigma^{\rm SI}_{\chi{\rm N}}$, $\sigma^{\rm LR}_{\rm\chi e}$ and $\sigma^{\rm SR}_{\rm\chi e}$) are all functions of $\rm m_{\chi}$. The input to the analysis, as depicted in Figure 1a, are the published AM amplitude data from these experiments: $\rm DL$ [14], COSINE [9] and ANAIS [15, 10] with NaI(Tl), XMASS with Xe [6] and CDEX with Ge [7]. These AM data sets are complementary in their strength in probing different parameter space. COSINE and ANAIS use the same detector technology with identical target isotopes as $\rm DL$. Comparisons would be model-independent in principle, though the variations on quenching factors among different NaI(Tl) crystals have been raised [20]. CDEX is distinguished with low detector threshold ($0.16~{}\rm keV_{ee}$) while XMASS has big exposure (1.82 ton-year). The electron-equivalent unit $\rm keV_{ee}$ is used to characterize detector response and measured energy ($\rm T$), unless otherwise stated.
The maximum(minimum) AM amplitude are on June 1(December 2) [21, 22] following the standard DM halo model [1, 4]. The quenching factors used by $\rm DL$ [23] is applied to all the three NaI(Tl) experiments. The expected AM differential spectra for positive WIMP signatures in $\rm\chi N^{SI}$ for the three targets are displayed in Figure 1b. The AM shapes originate from the differences in $\chi$-velocity relative to Earth between summer and winter. A characteristic feature is a drop from enhancement to deficit at low energy. The turning point is $\rm m_{\chi}$-dependent.
Figure 3: $\rm DL$ data [14] with best-fits under (a) ${\Gamma}_{\rm N}^{0}$: with $\chi{\rm N}$ only where $\rm m_{\chi}{=}54~{}{\rm GeV}$, and (b) ${\Gamma}_{tot}^{1\chi}$: with all three channels ($\rm\chi N^{SI}{,}\rm\chi e^{LR}{,}\rm\chi e^{SR}$) combined, in which case $\rm m_{\chi}{=}83~{}{\rm GeV}$.
Recent interest in the searches of light dark matter brings along intense activities in exploring novel detector concepts. Crucial to this development is the $\rm\chi e$ detection channels which triggers intense research efforts in the refinement on the evaluation of its interactions with atoms $A$ in $\chi{+}A{\rightarrow}\chi{+}e^{-}{+}A^{+}$ [16, 24]. Following our previous work on Ge and Xe with the many body approach with conventional frozen-core approximation (FCA) [16, 25, 26], we use same approach for Na and I prescribed in Ref. [16] in this analysis. The results are validated by comparing the derived photo-ionization cross-sections against experimental data [17] and previous calculations [18, 19]. The photo-ionization cross-sections of Na and I are illustrated in Figures 2a&b, showing consistency of theory calculations with measured data to within 5% across the entire energy transfer range. Identical results are obtained whether the targets are treated as atoms or ions. The differential AM spectra in $\rm\chi e^{LR}$ and $\rm\chi e^{SR}$ for the three target nuclei are shown in Figure 1c. The $\rm\chi e$ recoil energy is shifted lower in winter, giving rise to rapidly rising AM spectra. The long-range $\chi{\rm e}$ interactions has an additional $1/q^{2}$ term where $q$ is the 3-momentum transfer [16]. The differential AM spectra therefore rise steeper at low recoil energy such that studies of $\rm\chi e^{LR}$ favor experiments with lower detection threshold. The AM spectra in $\rm\chi e^{LR}$ and $\rm\chi e^{SR}$ for NaI are superimposed to the $\rm\chi N^{SI}$ spectra in Figure 1b, showing their vastly different response at low energy.
The best-fit estimators of the various cross-sections are derived by a minimum-$\chi^{2}$ analysis. At a given $\rm m_{\chi}$,
$\chi^{2}$ ${=}{\displaystyle\sum_{i}~{}\frac{1}{\Delta_{i}^{2}}}$$\Big{\{}n_{i}-\Big{[}{\sigma^{\rm SI}_{\chi{\rm N}}}~{}\phi_{\chi N}^{\rm SI}(\rm T_{i})+{\sigma^{\rm LR}_{\rm\chi e}}~{}\phi_{\rm\chi e}^{\rm LR}(\rm T_{i})$
$+{\sigma^{\rm SR}_{\rm\chi e}}~{}\phi_{\rm\chi e}^{\rm SR}(\rm T_{i})+{2\sqrt{\sigma^{\rm LR}_{\rm\chi e}\sigma^{\rm SR}_{\rm\chi e}}}~{}\phi_{\rm\chi e}^{\rm int}(\rm T_{i})\Big{]}\Big{\}}^{2}$
where $n_{i}$ and $\Delta_{i}$ are the AM amplitudes and uncertainties at the $i^{th}$-bin of average energy $\rm T_{i}$, while $\phi_{\rm\chi e}^{\rm SI}$, $\phi_{\rm\chi e}^{\rm LR}$, $\phi_{\rm\chi e}^{\rm SR}$ are the normalized spectral functions for the three interactions. The interference spectrum ($\phi_{\rm\chi e}^{\rm int}$) between $\rm\chi e^{LR}$ and $\rm\chi e^{SR}$ follows the many-body $\chi$-atom calculations given in Eq. 3 of Ref. [16], also depicted in Figure 1c.
The $\rm DL$ data shows positive AM signatures and reject the null hypothesis with large statistical significance. Its best-fit spectra with $\chi{\rm N}$-channel only (${\Gamma}_{\rm N}^{0}{:}~{}{\rm setting}~{}\sigma^{\rm LR}_{\rm\chi e}{=}\sigma^{\rm SR}_{\rm\chi e}{=}0$ in Eq. Dark Matter Annual Modulation Analysis with Combined Nuclear and Electron Recoil Channels) is displayed in Figure 3a. This analysis expands to have all three channels taken as the free fitting variables. Two DM scenarios are considered, parametrized by $f_{\chi}$ as the DM relic density fraction from the $\chi$ interacting via $\rm\chi e$: (1) ${\Gamma}_{tot}^{1\chi}-$ Both $\chi{\rm N}$ and $\rm\chi e$ interactions are due to a single $\chi$ ($f_{\chi}{=}1$), such that same constraints on $\rm m_{\chi}$ apply to all channels; (2) ${\Gamma}_{tot}^{2\chi}-$ The case of independent constraints where two different $\chi$’s with fractional density $f_{\chi}$ and $(1{-}f_{\chi})$ interact separately via $\rm\chi e$ and $\chi{\rm N}$, respectively. The limiting case of $f_{\chi}{=}0$ corresponds to the baseline ${\Gamma}_{\rm N}^{0}$. The analysis of ${\Gamma}_{tot}^{2\chi}$ is applicable to generic interactions by two independent WIMPs.
The combined best-fit in ${\Gamma}_{tot}^{1\chi}$ is given in Figure 3b. The allowed bands of $\sigma^{\rm SI}_{\chi{\rm N}}$ are derived with the Wilks’ Approximation [27]. It can be seen that the addition of the $\rm\chi e$-channels provides better description of the AM data at low energy near threshold, and the ${\Gamma}_{tot}^{1\chi}$-fit is a more inclusive and general form than the conventional ${\Gamma}_{\rm N}^{0}$. As illustrated in Table 1, the analysis reveal that the interpretation of the $\rm DL$ data incorporating ${\Gamma}_{tot}^{1\chi}$ or ${\Gamma}_{tot}^{2\chi}$ has higher statistical significance than ${\Gamma}_{\rm N}^{0}$ alone. The low energy data ($1{-}4~{}\rm keV_{ee}$) gives a p-values of 0.52 for ${\Gamma}_{tot}^{(1\chi,2\chi)}$ but only 0.07 for ${\Gamma}_{\rm N}^{0}$, indicating that the $\rm\chi N^{SI}$ channel as the sole physics scenario cannot explain the low energy data. In addition, the differences in $\chi^{2}/{\rm dof}$ between ${\Gamma}_{\rm N}^{0}$ and ${\Gamma}_{tot}^{(1\chi,2\chi)}$ (7.26/2 for ${\rm T}{=}1{-}4~{}\rm keV_{ee}$ and 9.66/2 for ${\rm T}{=}1{-}20~{}\rm keV_{ee}$) correspond to p-values of 0.02 and 0.008, respectively. This implies the necessity of having additional physical processes such as $\rm\chi e^{LR}$ and $\rm\chi e^{SR}$ to explain the AM spectrum [28].
${\Gamma}_{\rm N}^{0}$: ${\Gamma}_{tot}^{(1\chi,2\chi)}$: Compare
$\rm\chi N^{SI}$ $\rm\chi N^{SI}{+}\rm\chi e^{LR}{+}\rm\chi e^{SR}$
Data $\chi^{2}/{\rm dof}$ $\Delta\chi^{2}/{\rm dof}$
${}~{}(\rm keV_{ee})~{}~{}$ (p-value)    (p-value)
1 - 20    32.06/36 22.40/34 9.66/2
(0.66) (0.94) (0.008)
1 - 4 8.6/4 1.3/2 7.26/2
(0.07) (0.52) (0.02)
Table 1: Comparison of statistical significance (as $\chi^{2}/{\rm dof}$ and p-values) in the ${\Gamma}_{\rm N}^{0}$ and ${\Gamma}_{tot}^{(1\chi,2\chi)}$ fits to the DL AM data [14]. The increased significance from 1 to 3 fitting parameters indicates the presence of additional physical processes.
Figure 4: Exclusion Plots with AM Data from Na+I, Ge and Xe experiments with (a) $\sigma^{\rm SI}_{\chi{\rm N}}$ (b) $\sigma^{\rm LR}_{\rm\chi e}$, and (c) $\sigma^{\rm SR}_{\rm\chi e}$ versus $\rm m_{\chi}$. Y-axes are presented in $[f_{\chi}\sigma]$ to accommodate both scenarios ${\Gamma}_{tot}^{1\chi}$ $(f_{\chi}{=}1)$ and ${\Gamma}_{tot}^{2\chi}$ $(0{<}f_{\chi}{<}1)$. The DL allowed regions [14] from ${\Gamma}_{tot}^{1\chi}$ at 90%(±${\pm}3\sigma$) C.I. are presented as dark(light) shaded areas. The case of ${\Gamma}_{\rm N}^{0}$ is included in (a) as comparison. Exclusion contours from the null AM experiments represent upper limits at 90% C.L. Red circles of (b,c) are 90% allowed regions from ${\Gamma}_{tot}^{1\chi}$ for ($\sigma^{\rm LR}_{\rm\chi e}$ at $\rm\chi e^{SR}{=}0$,$\sigma^{\rm SR}_{\rm\chi e}$ at $\sigma^{\rm LR}_{\rm\chi e}{=}0$) while the blue bands correspond to those from ${\Gamma}_{tot}^{2\chi}$.
Figure 5: Correlations between $\sigma^{\rm SR}_{\rm\chi e}$ and $\sigma^{\rm LR}_{\rm\chi e}$, showing the allowed region in the 90% C.I. from $\rm DL$ AM data [14] and in comparison with the null AM experiments at upper limits at 90% C.L. (a) with the ${\Gamma}_{tot}^{1\chi}$ analysis, where the best-fit $\rm m_{\chi}$-range is $(60{\mbox{-}}100)~{}{\rm GeV}$, and with the ${\Gamma}_{tot}^{2\chi}$ analysis, at (b) $\rm m_{\chi}{=}(100{\pm}10)~{}{\rm GeV}$, and (c) $\rm m_{\chi}{=}(1{\pm}0.1)~{}{\rm GeV}$.
On the contrary, no AM signatures were observed in the COSINE, ANAIS, XMASS, CDEX experiments. Their data are consistent with the null hypothesis. Conservative and inclusive limits are derived with single-channel (${\Gamma}_{\rm N}^{0}$) analysis -- $\sigma^{\rm LR}_{\rm\chi e}$ and $\sigma^{\rm SR}_{\rm\chi e}$ are set to be zero in Eq. Dark Matter Annual Modulation Analysis with Combined Nuclear and Electron Recoil Channels while the best-fit values of $\sigma^{\rm SI}_{\chi{\rm N}}$ are evaluated, allowing negative values. Limits are derived with the unified approach [29]. Same procedures apply for $\sigma^{\rm LR}_{\rm\chi e}$ and $\sigma^{\rm SR}_{\rm\chi e}$ by assuming two other $\sigma$’s are zero.
The allowed regions in both ±${\pm}3\sigma$ and 90% confidence intervals (C.I.) for $\rm\chi N^{SI}$, $\rm\chi e^{LR}$ and $\rm\chi e^{SR}$ as functions of $\rm m_{\chi}$ under ${\Gamma}_{tot}^{1\chi}$-analysis from $\rm DL$ AM data [14] are shown, respectively, in Figures 4a,b&c. The exclusion contours showing upper limits at 90% confidence level (C.L.) from the null AM experiments are superimposed for comparisons in the same plots. Results of ${\Gamma}_{\rm N}^{0}$ are superimposed in Figure 4a for comparison. The relevant ranges of the Earth attenuation and scattering effects correspond to much higher cross-sections and the bounds are outside to the plots displayed. For instance for the CDEX experiment at 2.4 km of rock overburden, the upper limits of the sensitivity regions are $10^{\mbox{-}30}~{}{\rm cm}^{2}$ for $\sigma^{\rm SI}_{\chi{\rm N}}$ [30], $10^{\mbox{-}24}~{}{\rm cm}^{2}$ for $\sigma^{\rm LR}_{\rm\chi e}$, and $10^{\mbox{-}29}~{}{\rm cm}^{2}$ for $\sigma^{\rm SR}_{\rm\chi e}$.
There are two allowed regions in ${\Gamma}_{\rm N}^{0}$ at 90% C.I. at low and high $\rm m_{\chi}$ corresponding to Na and I-recoils, respectively. With $\rm\chi e^{LR}$ and $\rm\chi e^{SR}$ are added in ${\Gamma}_{tot}^{1\chi}$, only the high $\rm m_{\chi}$ region due to I-recoils is allowed at the same significance. The best-fit solution of $\rm m_{\chi}$ is shifted from $(54.25{\pm}4.25)~{}{\rm GeV}$ in ${\Gamma}_{\rm N}^{0}$ to $(83.3{\pm}25.65)~{}{\rm GeV}$ in ${\Gamma}_{tot}^{1\chi}$. The $\rm\chi e$-channels dominate the near-threshold behavior, while the high energy (${>}3~{}\rm keV_{ee}$) spectra are defined by $\chi{\rm N}$. The two $\rm\chi e$ channels are highly correlated and share the strength of the AM counts at low energy. The limiting cases of the best-fit values in ${\Gamma}_{tot}^{1\chi}$ of $\sigma^{\rm LR}_{\rm\chi e}$ at $\sigma^{\rm SR}_{\rm\chi e}{=}0$ and $\sigma^{\rm SR}_{\rm\chi e}$ at $\sigma^{\rm LR}_{\rm\chi e}{=}0$ correspond to the red circles in Figures 4b&c, respectively. The exclusion contours of the null AM experiments are superimposed. The weaker bounds in $\sigma^{\rm LR}_{\rm\chi e}$ in Figure 4b originates from its sharper rise of the AM spectra due to an additional $(1/q^{2})$ dependence, so that only the low threshold data from CDEX places strong constraints. Under ${\Gamma}_{tot}^{1\chi}$, the DL AM allowed regions in Figures 4a,b&c are inter-dependent -- $\rm m_{\chi}$ is constrained to $(60{\mbox{-}}100)~{}{\rm GeV}$ by the $\rm\chi N^{SI}$ channel. The correlations of $(\sigma^{\rm LR}_{\rm\chi e},\sigma^{\rm SR}_{\rm\chi e})$ in this $\rm m_{\chi}$ range are depicted in Figure 5a, showing the DL allowed region in ${\Gamma}_{tot}^{1\chi}$ is excluded by the null AM experiments, and in particular with large margins by CDEX and XMASS combined.
In the scenario of ${\Gamma}_{tot}^{2\chi}$ where $\chi{\rm N}$ and $\rm\chi e$ interactions are due to two different $\chi$’s, the constraints due to the high and low energy AM spectral components are independent. The $\rm DL$ $\rm\chi N^{SI}$ allowed regions are defined by the high-energy component and remain those of Figure 4a, which are well-excluded by the other null experiments. The AM low-energy component is dominated by the $\rm\chi e$ channels of a different $\chi$ unconstrained in $\rm m_{\chi}$. This gives rise to allowed regions in ${\Gamma}_{tot}^{2\chi}$ represented by the blue bands in Figures 4b&c for $\sigma^{\rm LR}_{\rm\chi e}$ and $\sigma^{\rm SR}_{\rm\chi e}$, respectively. As illustration, the $\rm DL$ 90% allowed C.I. for $\rm m_{\chi}{=}(100{\pm}10)~{}{\rm and}~{}(1{\pm}0.1)~{}{\rm GeV}$ are depicted in Figure 5b&c, respectively, superimposed with the exclusion contours of the null AM data. It can be concluded that in the ${\Gamma}_{tot}^{2\chi}$ analysis, the DL AM allowed regions associated with $\rm\chi e$ are probed and rejected by the various null AM experiments. The CDEX [7] data placed more severe constraints on $\sigma^{\rm LR}_{\rm\chi e}$ where the $(1/q^{2})$ dependence favors low threshold experiments. The XMASS [6] experiment is more sensitive to $\sigma^{\rm SR}_{\rm\chi e}$ due to its large exposure. It can be projected from the results that the DL allowed regions in generic two-WIMP models are ruled out by the null AM experiments. Very little room is left to account for the DL AM data with WIMP-induced effects.
This work is supported by contracts 106-2923-M-001-006-MY5, 110-2112-M-001-029-MY3, and 113-2112-M-001-053-MY3 from the National Science and Technology Council, Taiwan, grants 2021-22/TG2.1 from the National Center of Theoretical Sciences, Taiwan.

### v2

@@
 It has been shown that $\rm\chi e$ (WIMP interactions with electrons) alone is not compatible with the DL AM data.
-We expand the investigations by performing an AM analysis with the addition of $\rm\chi e$ long-range and short-range interactions to $\chi{\rm N}$, derived using the frozen-core approximation method.
+We expand the investigations by performing an AM analysis with the addition of $\rm\chi e$ long-range and short-range interactions to $\chi{\rm N}$, derived using the Frozen Core Approximation method.
 Two scenarios are considered, where the $\chi{\rm N}$ and $\rm\chi e$ processes are due to a single $\chi$ (${\Gamma}_{tot}^{1\chi}$) or two different $\chi$’s (${\Gamma}_{tot}^{2\chi}$).
@@
 It can be projected that DL-allowed parameter spaces from generic models with interactions induced by two-WIMPs are ruled out.
+I Introduction
 There are compelling experimental evidence that about one-quarter of the energy density of the Universe is composed of dark matter (DM), whose exact nature and properties remain unknown.
@@
 Direct experimental searches assume finite interactions between WIMPs with electrons ($\rm\chi e$) and nuclei ($\chi{\rm N}$).
-Positive signatures manifest as excess events over known background in the measured “time-integrated” (TI) energy spectra and in their annual modulation (AM) due to changes of the relative velocity between Earth and the WIMPs in the galactic halo [4].
+Positive signatures manifest as excess events over known background in the measured “time-integrated” (TI) energy spectra and in their annual modulation (AM) due to changes of the relative velocity between the Earth and the WIMPs in the galactic halo [4].
 TI analysis is sensitive to uncertainties of background modeling while AM analysis only requires the background is stable with time but independent of other details.
@@
 This interpretation, however, is challenged and rejected by numerous $\chi{\rm N}$ experiments with TI analysis [1] and several with AM analysis [6, 7, 8, 9, 10] using a variety of targets.
+In particular, the COSINE [9] and ANAIS [10] experiments adopt the same NaI(Tl) as target nuclei as DL.
 There are attempts to explain the $\rm DL$ AM data with scenarios other than $\chi{\rm N}$ detection, such as complications in the analysis procedures [11, 12].
-In addition, it has been shown [13] -- and verified in our studies -- that AM analysis with the $\rm\chi e$-channels only (setting $\sigma^{\rm SI}_{\chi{\rm N}}{=}0$ in Eq.
-Dark Matter Annual Modulation Analysis with Combined Nuclear and Electron Recoil Channels) is unable to provide an acceptable fit to the $\rm DL$ data.
-The maximum electron recoil energy even for infinite $\rm m_{\chi}$ is about $3~{}\rm keV_{ee}$, in serious discrepancy with the data.
+Table 1: The AM experiments whose data are selected for this analysis. Their relevant experimental configurations are listed.
+Experiment Data Set Detector      Mass Live Time/Duration    Exposure Threshold Ref.
+   (kg) (year) (ton-year) (keVee)
+DAMA/LIBRA DAMA NaI (Tl) 87.3 3.32/6.85 0.29 2.0
+     (DL) DL-Phase 1 242.5 4.29/7.10 1.04 2.0 [13]
+DL-Phase 2 242.5 4.66/5.89 1.13 1.0
+COSINE    COSINE-100 NaI (Tl) 61.3 5.80/6.40 0.358 0.7 [9]
+ANAIS ANAIS-112 NaI (Tl) 112.5 2.87/3.03 0.323 1.0 [14, 10]
+CDEX CDEX-1B Ge 0.939 3.20/4.20 0.00303 0.25 [7]
+XMASS XMASS-I Liquid Xe 832.0 2.19/2.70 1.82 1.0 [6]
+As elaborated in Section III and summarized in Table 2, there exists tension to interpret the DL data at ($1{-}4~{}\rm keV_{ee}$) exclusively by a $\chi{\rm N}$-only scenario.
+The p-value of 0.07 implies additional effects are preferred that contribute to the DL AM low energy data.
+Analogously, it has been demonstrated [15] -- and verified in our studies -- that AM analysis with the $\rm\chi e$-only channel is unable to provide an acceptable fit to the $\rm DL$ data.
+The AM rates are negligible above $3~{}\rm keV_{ee}$ in relevant range for this work of $\rm m_{\chi}{<}10^{3}~{}{\rm GeV}$, in serious discrepancy with the data.
+In comparison, when a $\chi{\rm N}{+}\rm\chi e$ scenario is considered in AM analysis, a p-value of 0.008 is obtained for its $\chi 2$-difference with the $\chi{\rm N}$-only case.
+This indicates that the addition of $\rm\chi e$ interactions to $\chi{\rm N}$ is statistically significant to match the DL low energy AM data.
+This article is organized as follows.
+We discuss the selection of input data in this analysis in Section II, followed by the evaluation of the $\rm\chi e$ cross sections and the statistical analysis procedures.
+The results and their interpretations are presented in Section III, in which the two cases of the $\chi{\rm N}$ and $\rm\chi e$ interactions originated from the same or different $\chi$s are treated separately.
 Figure 1: (a) Published AM amplitudes data from DM experiments with NaI in $\rm DL$ [13], COSINE [9] and ANAIS [14, 10], Xe in XMASS [6] and Ge in CDEX [7].
-For display purposes, the depicted event rates are re-scaled to account for the differences in bin-sizes and target mass number squared.
+To allow the different nuclei be presented with a common axis for display purposes, the depicted CDEX event rates are re-scaled by a factor 0.04.
 The expected differential AM spectra for positive WIMP signatures in: (b) $\rm\chi N^{SI}$ and (c) in $\rm\chi e^{LR}$ and $\rm\chi e^{SR}$ for all three targets taking the case of $\rm m_{\chi}{=}80~{}{\rm GeV}$ as example, together with the interference spectrum between $\rm\chi e^{LR}$ and $\rm\chi e^{SR}$ for NaI.
-The $\rm\chi e^{LR}$ and $\rm\chi e^{SR}$ spectra in (c) are superimposed as dotted lines in (b) f showing their responses to AM are very different from $\rm\chi N^{SI}$.
-The interference spectrum between $\rm\chi e^{LR}$ and $\rm\chi e^{SR}$ for NaI is also displayed in (c).
+The $\rm\chi e^{LR}$ and $\rm\chi e^{SR}$ spectra in (c) are superimposed as dotted lines in (b) showing their responses to AM are very different from $\rm\chi N^{SI}$.
 Figure 2: Photo-ionization cross-sections derived by FCA following Ref.
-[16] for (a) Na and (b) I, both as atoms and ions, within 1 $\rm keV_{ee}$ to 30 $\rm keV_{ee}$ energy transfer, The experimental data [17] and theoretical values [18, 19] are superimposed.
+[16] for (a) Na and (b) I, both as atoms and ions, within 1 $\rm keV_{ee}$ to 30 $\rm keV_{ee}$ energy transfer.
+The experimental data [17] and theoretical values [18, 19] are superimposed.
 Relative deviations from benchmark data are displayed.
+II Formulation
 We performed an analysis including both $\rm\chi e$ and $\chi{\rm N}$ interactions.
-The combined analysis would provide better matching statistically to the $\rm DL$ data.
+The combined analysis would provide improved statistical fit to the $\rm DL$ data.
 Three types of WIMPs interactions are included: spin-independent WIMP-nucleon interaction ($\rm\chi N^{SI}$), long and short-range WIMP-electron interaction ($\rm\chi e^{LR}$ and $\rm\chi e^{SR}$, respectively) [16].
 Their cross-sections (respectively $\sigma^{\rm SI}_{\chi{\rm N}}$, $\sigma^{\rm LR}_{\rm\chi e}$ and $\sigma^{\rm SR}_{\rm\chi e}$) are all functions of $\rm m_{\chi}$.
-The input to the analysis, as depicted in Figure 1a, are the published AM amplitude data from these experiments: $\rm DL$ [14], COSINE [9] and ANAIS [15, 10] with NaI(Tl), XMASS with Xe [6] and CDEX with Ge [7].
+II.1 Input Data
+The input to the analysis, as depicted in Figure 1a, are the published AM amplitudes data.
+A summary of the relevant features of the AM experiments is given in Table 1.
 These AM data sets are complementary in their strength in probing different parameter space.
-COSINE and ANAIS use the same detector technology with identical target isotopes as $\rm DL$.
+The $\rm DL$ [13] experiment is the only one reporting positive AM results.
+COSINE [9] and ANAIS [14, 10] use the same detector technology -- NaI(Tl) crystal scintillators -- with identical target isotopes as $\rm DL$.
 Comparisons would be model-independent in principle, though the variations on quenching factors among different NaI(Tl) crystals have been raised [20].
-CDEX is distinguished with low detector threshold ($0.16~{}\rm keV_{ee}$) while XMASS has big exposure (1.82 ton-year).
+CDEX is distinguished with low detector threshold ($0.25~{}\rm keV_{ee}$ for the AM analysis) made possible by novel p-type point-contact germanium detectors, while XMASS is a single-phase liquid xenon detector with large exposure (1.82 ton-year).
 The electron-equivalent unit $\rm keV_{ee}$ is used to characterize detector response and measured energy ($\rm T$), unless otherwise stated.
-The maximum(minimum) AM amplitude are on June 1(December 2) [21, 22] following the standard DM halo model [1, 4].
+The maximum(minimum) AM amplitudes are on June 1(December 2) [21, 22] following the standard DM halo model [1, 4].
 The quenching factors used by $\rm DL$ [23] is applied to all the three NaI(Tl) experiments.
 The expected AM differential spectra for positive WIMP signatures in $\rm\chi N^{SI}$ for the three targets are displayed in Figure 1b.
-The AM shapes originate from the differences in $\chi$-velocity relative to Earth between summer and winter.
+The AM spectral shapes originate from the differences in $\chi$-velocity relative to the Earth between summer and winter.
 A characteristic feature is a drop from enhancement to deficit at low energy.
 The turning point is $\rm m_{\chi}$-dependent.
-Figure 3: $\rm DL$ data [14] with best-fits under (a) ${\Gamma}_{\rm N}^{0}$: with $\chi{\rm N}$ only where $\rm m_{\chi}{=}54~{}{\rm GeV}$, and (b) ${\Gamma}_{tot}^{1\chi}$: with all three channels ($\rm\chi N^{SI}{,}\rm\chi e^{LR}{,}\rm\chi e^{SR}$) combined, in which case $\rm m_{\chi}{=}83~{}{\rm GeV}$.
-Recent interest in the searches of light dark matter brings along intense activities in exploring novel detector concepts.
-Crucial to this development is the $\rm\chi e$ detection channels which triggers intense research efforts in the refinement on the evaluation of its interactions with atoms $A$ in $\chi{+}A{\rightarrow}\chi{+}e^{-}{+}A^{+}$ [16, 24].
-Following our previous work on Ge and Xe with the many body approach with conventional frozen-core approximation (FCA) [16, 25, 26], we use same approach for Na and I prescribed in Ref.
-[16] in this analysis.
+II.2 Cross-Sections in $\rm\chi e$
+Recent interest in the searches of light DM brings along intense activities in exploring novel detector concepts.
+Crucial to this development is the $\rm\chi e$ detection channels which triggers intense research efforts in the refinement on the evaluation of its interactions with atoms $A$ [16, 24] in
+$\chi{+}A{\rightarrow}\chi{+}e^{-}{+}A^{+}~{}~{}.$ (1)
+Frozen Core Approximation (FCA) [25, 26], is a well-established framework in atomic many-body physics, recognized for its effectiveness in unraveling complex interactions and dynamics within atomic systems.
+Following our previous work on Ge and Xe with FCA [16, 24], we extend the same approach to Na and I in this work.
 The results are validated by comparing the derived photo-ionization cross-sections against experimental data [17] and previous calculations [18, 19].
-The photo-ionization cross-sections of Na and I are illustrated in Figures 2a&b, showing consistency of theory calculations with measured data to within 5% across the entire energy transfer range.
+The photo-ionization cross-sections of Na and I are illustrated in Figures 2a&b, showing consistency of theory calculations with measured data to within 5% across the relevant energy transfer range of $1{-}30~{}\rm keV_{ee}$.
 Identical results are obtained whether the targets are treated as atoms or ions.
-The differential AM spectra in $\rm\chi e^{LR}$ and $\rm\chi e^{SR}$ for the three target nuclei are shown in Figure 1c.
+The consistency indicates FCA can provide reliable modeling to the interactions of $\chi$ with the atoms.
+The differential AM spectra in $\rm\chi e^{LR}$ and $\rm\chi e^{SR}$ for the three targets are shown in Figure 1c.
 The $\rm\chi e$ recoil energy is shifted lower in winter, giving rise to rapidly rising AM spectra.
+In practice, only data ${<}4~{}\rm keV_{ee}$ would contribute to the analysis of the $\rm\chi e$ channels.
 The long-range $\chi{\rm e}$ interactions has an additional $1/q^{2}$ term where $q$ is the 3-momentum transfer [16].
 The differential AM spectra therefore rise steeper at low recoil energy such that studies of $\rm\chi e^{LR}$ favor experiments with lower detection threshold.
-The AM spectra in $\rm\chi e^{LR}$ and $\rm\chi e^{SR}$ for NaI are superimposed to the $\rm\chi N^{SI}$ spectra in Figure 1b, showing their vastly different response at low energy.
+The AM spectra in $\rm\chi e^{LR}$ and $\rm\chi e^{SR}$ for the three targets are superimposed to the $\rm\chi N^{SI}$ spectra in Figure 1b, showing their vastly different response at low energy.
+Figure 3: $\rm DL$ data [13] with best-fits under (a) ${\Gamma}_{\rm N}^{0}$: with $\chi{\rm N}$ only where $\rm m_{\chi}{=}54~{}{\rm GeV}$, and (b) ${\Gamma}_{tot}^{(1\chi,2\chi)}$: with all three channels ($\rm\chi N^{SI}{,}\rm\chi e^{LR}{,}\rm\chi e^{SR}$) combined, in which case $\rm m_{\chi}{=}83~{}{\rm GeV}$.
+II.3 Analysis
 The best-fit estimators of the various cross-sections are derived by a minimum-$\chi^{2}$ analysis. At a given $\rm m_{\chi}$,
@@
 Its best-fit spectra with $\chi{\rm N}$-channel only (${\Gamma}_{\rm N}^{0}{:}~{}{\rm setting}~{}\sigma^{\rm LR}_{\rm\chi e}{=}\sigma^{\rm SR}_{\rm\chi e}{=}0$ in Eq.
-Dark Matter Annual Modulation Analysis with Combined Nuclear and Electron Recoil Channels) is displayed in Figure 3a.
+II.3) is displayed in Figure 3a.
 This analysis expands to have all three channels taken as the free fitting variables.
-Two DM scenarios are considered, parametrized by $f_{\chi}$ as the DM relic density fraction from the $\chi$ interacting via $\rm\chi e$: (1) ${\Gamma}_{tot}^{1\chi}-$ Both $\chi{\rm N}$ and $\rm\chi e$ interactions are due to a single $\chi$ ($f_{\chi}{=}1$), such that same constraints on $\rm m_{\chi}$ apply to all channels; (2) ${\Gamma}_{tot}^{2\chi}-$ The case of independent constraints where two different $\chi$’s with fractional density $f_{\chi}$ and $(1{-}f_{\chi})$ interact separately via $\rm\chi e$ and $\chi{\rm N}$, respectively.
+Two DM scenarios are considered, parametrized by $f_{\chi}$ as the DM relic density fraction from the $\chi$ interacting via $\rm\chi e$:
+  1. 1.
+    ${\Gamma}_{tot}^{1\chi}-$ Both $\chi{\rm N}$ and $\rm\chi e$ interactions are due to a single $\chi$ ($f_{\chi}{=}1$), such that same constraints on $\rm m_{\chi}$ apply to all channels;
+  2. 2.
+    ${\Gamma}_{tot}^{2\chi}-$ The case of independent constraints where two different $\chi$’s with fractional density $f_{\chi}$ and $(1{-}f_{\chi})$ interact separately via $\rm\chi e$ and $\chi{\rm N}$, respectively.
 The limiting case of $f_{\chi}{=}0$ corresponds to the baseline ${\Gamma}_{\rm N}^{0}$.
-The analysis of ${\Gamma}_{tot}^{2\chi}$ is applicable to generic interactions by two independent WIMPs.
-The combined best-fit in ${\Gamma}_{tot}^{1\chi}$ is given in Figure 3b.
-The allowed bands of $\sigma^{\rm SI}_{\chi{\rm N}}$ are derived with the Wilks’ Approximation [27].
-It can be seen that the addition of the $\rm\chi e$-channels provides better description of the AM data at low energy near threshold, and the ${\Gamma}_{tot}^{1\chi}$-fit is a more inclusive and general form than the conventional ${\Gamma}_{\rm N}^{0}$.
-As illustrated in Table 1, the analysis reveal that the interpretation of the $\rm DL$ data incorporating ${\Gamma}_{tot}^{1\chi}$ or ${\Gamma}_{tot}^{2\chi}$ has higher statistical significance than ${\Gamma}_{\rm N}^{0}$ alone.
-The low energy data ($1{-}4~{}\rm keV_{ee}$) gives a p-values of 0.52 for ${\Gamma}_{tot}^{(1\chi,2\chi)}$ but only 0.07 for ${\Gamma}_{\rm N}^{0}$, indicating that the $\rm\chi N^{SI}$ channel as the sole physics scenario cannot explain the low energy data.
-In addition, the differences in $\chi^{2}/{\rm dof}$ between ${\Gamma}_{\rm N}^{0}$ and ${\Gamma}_{tot}^{(1\chi,2\chi)}$ (7.26/2 for ${\rm T}{=}1{-}4~{}\rm keV_{ee}$ and 9.66/2 for ${\rm T}{=}1{-}20~{}\rm keV_{ee}$) correspond to p-values of 0.02 and 0.008, respectively.
-This implies the necessity of having additional physical processes such as $\rm\chi e^{LR}$ and $\rm\chi e^{SR}$ to explain the AM spectrum [28].
+Table 2: Comparison of the statistical significance in terms of $\chi^{2}/{\rm dof}$ of the ${\Gamma}_{\rm N}^{0}$ and ${\Gamma}_{tot}^{(1\chi,2\chi)}$ fitted to the published DL AM data [13].
+The p-values which qualify likelihood of tested hypothesis are also shown.
+The results can be compared to those of a test case with 20% reduced uncertainties which produces p=0.5 on the complete data set (*).
+The increased significance from ${\Gamma}_{\rm N}^{0}$ to ${\Gamma}_{tot}^{(1\chi,2\chi)}$ implies the presence of genuine physical processes.
 ${\Gamma}_{\rm N}^{0}$: ${\Gamma}_{tot}^{(1\chi,2\chi)}$: Compare
@@
 ${}~{}(\rm keV_{ee})~{}~{}$ (p-value)    (p-value)
+Published Data
 1 - 20    32.06/36 22.40/34 9.66/2
@@
 (0.07) (0.52) (0.02)
-Table 1: Comparison of statistical significance (as $\chi^{2}/{\rm dof}$ and p-values) in the ${\Gamma}_{\rm N}^{0}$ and ${\Gamma}_{tot}^{(1\chi,2\chi)}$ fits to the DL AM data [14].
-The increased significance from 1 to 3 fitting parameters indicates the presence of additional physical processes.
+Test Case (Effects of Reduced Uncertainties)*
+1 - 20    48.1/36 33.33/34 14.8/2
+(0.086) (0.50)* (0.0006)
+1 - 4 12.8/4 1.94/2 10.86/2
+(0.012) (0.38) (0.0044)
+Figure 4: Predicted spectra for the null AM experiments: (a) COSINE, ANAIS, XMASS, and (b) CDEX, due to the ${\Gamma}_{tot}^{(1\chi,2\chi)}$ best-fit values of ($\rm\chi N^{SI}{,}\rm\chi e^{LR}{,}\rm\chi e^{SR}$) on the DL AM data in Figure 3.
 Figure 5: Exclusion plots with AM Data from Na+I, Ge and Xe experiments with (a) $\sigma^{\rm SI}_{\chi{\rm N}}$ (b) $\sigma^{\rm LR}_{\rm\chi e}$, and (c) $\sigma^{\rm SR}_{\rm\chi e}$ versus $\rm m_{\chi}$.
@@
 Exclusion contours from the null AM experiments represent upper limits at 90% C.L.
-Red circles of (b,c) are 90% allowed regions from ${\Gamma}_{tot}^{1\chi}$ for ($\sigma^{\rm LR}_{\rm\chi e}$ at $\rm\chi e^{SR}{=}0$,$\sigma^{\rm SR}_{\rm\chi e}$ at $\sigma^{\rm LR}_{\rm\chi e}{=}0$) while the blue bands correspond to those from ${\Gamma}_{tot}^{2\chi}$.
+The light red circles of (b,c) are allowed regions at 90% C.I.
+from ${\Gamma}_{tot}^{1\chi}$ for ($\sigma^{\rm LR}_{\rm\chi e}$ at $\sigma^{\rm SR}_{\rm\chi e}{=}0$,$\sigma^{\rm SR}_{\rm\chi e}$ at $\sigma^{\rm LR}_{\rm\chi e}{=}0$) while the light blue bands correspond to those from ${\Gamma}_{tot}^{2\chi}$.
+III Results and Interpretations
+The combined best-fit in ${\Gamma}_{tot}^{1\chi}$ is given in Figure 3b.
+The spectra for ${\Gamma}_{tot}^{2\chi}$ would be identical.
+The allowed bands of $\sigma^{\rm SI}_{\chi{\rm N}}$ are derived with the Wilks’ approximation [27].
+The addition of the $\rm\chi e$-channels provides better description of the DL AM data ${<}4~{}\rm keV_{ee}$.
+The ${\Gamma}_{tot}^{(1\chi,2\chi)}$-fit is a more inclusive and general form than the conventional ${\Gamma}_{\rm N}^{0}$.
+As illustrated in Table 2, the analysis reveal that the interpretation of the $\rm DL$ published data [13] incorporating ${\Gamma}_{tot}^{(1\chi,2\chi)}$ has higher statistical significance than ${\Gamma}_{\rm N}^{0}$ alone.
+The low energy data ($1{-}4~{}\rm keV_{ee}$) gives p-values of 0.52 for ${\Gamma}_{tot}^{(1\chi,2\chi)}$ but only 0.07 for ${\Gamma}_{\rm N}^{0}$, indicating that the $\rm\chi N^{SI}$ channel as the sole physics scenario cannot explain the low energy data.
+In addition, the differences in $\chi^{2}/{\rm dof}$ between ${\Gamma}_{\rm N}^{0}$ and ${\Gamma}_{tot}^{(1\chi,2\chi)}$ (7.26/2 for ${\rm T}{=}1{-}4~{}\rm keV_{ee}$ and 9.66/2 for ${\rm T}{=}1{-}20~{}\rm keV_{ee}$) correspond to p-values of 0.02 and 0.008, respectively.
+This implies the necessity of having additional physical processes such as $\rm\chi e^{LR}$ and $\rm\chi e^{SR}$ to explain the AM spectrum [28].
+We note the $\chi^{2}/{\rm dof}$ value for the complete 1--20 keV data set in ${\Gamma}_{tot}^{(1\chi,2\chi)}$ in Table 2 is significantly less than 1.
+This suggests a scenario where the published uncertainties [13] are over-estimated.
+We investigate a test case where the uncertainties of the DL data are uniformly reduced by 20%, resulting in p=0.5.
+The changes to $\chi^{2}/{\rm dof}$ and p-values are shown in Table 2.
+The tension against ${\Gamma}_{\rm N}^{0}$ as a valid hypothesis is stronger while ${\Gamma}_{tot}^{(1\chi,2\chi)}$ shows perfect agreement with data.
 Figure 6: Correlations between $\sigma^{\rm SR}_{\rm\chi e}$ and $\sigma^{\rm LR}_{\rm\chi e}$, showing the allowed region in the 90% C.I.
-from $\rm DL$ AM data [14] and in comparison with the null AM experiments at upper limits at 90% C.L.
+from $\rm DL$ AM data [13] and in comparison with the null AM experiments at 90% C.L.
+upper limits.
 (a) with the ${\Gamma}_{tot}^{1\chi}$ analysis, where the best-fit $\rm m_{\chi}$-range is $(60{\mbox{-}}100)~{}{\rm GeV}$, and with the ${\Gamma}_{tot}^{2\chi}$ analysis, at (b) $\rm m_{\chi}{=}(100{\pm}10)~{}{\rm GeV}$, and (c) $\rm m_{\chi}{=}(1{\pm}0.1)~{}{\rm GeV}$.
@@
 Their data are consistent with the null hypothesis.
+As illustrations, the predicted AM spectra in these experiments due to the best-fit values of ($\rm\chi N^{SI}{,}\rm\chi e^{LR}{,}\rm\chi e^{SR}$) derived from DL data are depicted in Figures 4a&b.
+The incompatibility of the DL best-fit values with the null AM experiments can be readily visualized.
 Conservative and inclusive limits are derived with single-channel (${\Gamma}_{\rm N}^{0}$) analysis -- $\sigma^{\rm LR}_{\rm\chi e}$ and $\sigma^{\rm SR}_{\rm\chi e}$ are set to be zero in Eq.
-Dark Matter Annual Modulation Analysis with Combined Nuclear and Electron Recoil Channels while the best-fit values of $\sigma^{\rm SI}_{\chi{\rm N}}$ are evaluated, allowing negative values.
+II.3 while the best-fit values of $\sigma^{\rm SI}_{\chi{\rm N}}$ are evaluated, allowing negative values.
 Limits are derived with the unified approach [29].
 Same procedures apply for $\sigma^{\rm LR}_{\rm\chi e}$ and $\sigma^{\rm SR}_{\rm\chi e}$ by assuming two other $\sigma$’s are zero.
+III.1 Case of ${\Gamma}_{tot}^{1\chi}$
 The allowed regions in both ±${\pm}3\sigma$ and 90% confidence intervals (C.I.) for $\rm\chi N^{SI}$, $\rm\chi e^{LR}$ and $\rm\chi e^{SR}$ as functions of $\rm m_{\chi}$ under ${\Gamma}_{tot}^{1\chi}$-analysis from $\rm DL$ AM data [13] are shown, respectively, in Figures 5a,b&c.
@@
 The two $\rm\chi e$ channels are highly correlated and share the strength of the AM counts at low energy.
-The limiting cases of the best-fit values in ${\Gamma}_{tot}^{1\chi}$ of $\sigma^{\rm LR}_{\rm\chi e}$ at $\sigma^{\rm SR}_{\rm\chi e}{=}0$ and $\sigma^{\rm SR}_{\rm\chi e}$ at $\sigma^{\rm LR}_{\rm\chi e}{=}0$ correspond to the red circles in Figures 4b&c, respectively.
+The limiting cases of the best-fit values in ${\Gamma}_{tot}^{1\chi}$ of $\sigma^{\rm LR}_{\rm\chi e}$ at $\sigma^{\rm SR}_{\rm\chi e}{=}0$ and $\sigma^{\rm SR}_{\rm\chi e}$ at $\sigma^{\rm LR}_{\rm\chi e}{=}0$ correspond to the light red circles in Figures 5b&c, respectively.
 The exclusion contours of the null AM experiments are superimposed.
@@
 Under ${\Gamma}_{tot}^{1\chi}$, the DL AM allowed regions in Figures 5a,b&c are inter-dependent -- $\rm m_{\chi}$ is constrained to $(60{\mbox{-}}100)~{}{\rm GeV}$ by the $\rm\chi N^{SI}$ channel.
-The correlations of $(\sigma^{\rm LR}_{\rm\chi e},\sigma^{\rm SR}_{\rm\chi e})$ in this $\rm m_{\chi}$ range are depicted in Figure 5a, showing the DL allowed region in ${\Gamma}_{tot}^{1\chi}$ is excluded by the null AM experiments, and in particular with large margins by CDEX and XMASS combined.
+Figures 5b(c) shows that the low cross-section portion of $\sigma^{\rm LR}_{\rm\chi e}$($\sigma^{\rm SR}_{\rm\chi e}$) allowed regions at this $\rm m_{\chi}$ range remains unprobed by its own interactions.
+However, this parameter space is correlated with the high cross-section portion of its counterparts.
+The correlations are depicted in Figure 6a which shows the combined constraints of $(\sigma^{\rm LR}_{\rm\chi e},\sigma^{\rm SR}_{\rm\chi e})$ from the null AM experiments can reject the entire DL allowed region in ${\Gamma}_{tot}^{1\chi}$.
+The exclusion margins are particularly large through the combined constraints of CDEX and XMASS.
+III.2 Case of ${\Gamma}_{tot}^{2\chi}$
 In the scenario of ${\Gamma}_{tot}^{2\chi}$ where $\chi{\rm N}$ and $\rm\chi e$ interactions are due to two different $\chi$’s, the constraints due to the high and low energy AM spectral components are independent.
@@
 The AM low-energy component is dominated by the $\rm\chi e$ channels of a different $\chi$ unconstrained in $\rm m_{\chi}$.
-This gives rise to allowed regions in ${\Gamma}_{tot}^{2\chi}$ represented by the blue bands in Figures 4b&c for $\sigma^{\rm LR}_{\rm\chi e}$ and $\sigma^{\rm SR}_{\rm\chi e}$, respectively.
-As illustration, the $\rm DL$ 90% allowed C.I.
-for $\rm m_{\chi}{=}(100{\pm}10)~{}{\rm and}~{}(1{\pm}0.1)~{}{\rm GeV}$ are depicted in Figure 5b&c, respectively, superimposed with the exclusion contours of the null AM data.
-It can be concluded that in the ${\Gamma}_{tot}^{2\chi}$ analysis, the DL AM allowed regions associated with $\rm\chi e$ are probed and rejected by the various null AM experiments.
+This gives rise to allowed regions in ${\Gamma}_{tot}^{2\chi}$ represented by the light blue bands in Figures 5b&c for $\sigma^{\rm LR}_{\rm\chi e}$ and $\sigma^{\rm SR}_{\rm\chi e}$, respectively.
+The ${\Gamma}_{tot}^{2\chi}$ scenario is excluded by the null AM experiments, independently in $\sigma^{\rm LR}_{\rm\chi e}$ and $\sigma^{\rm SR}_{\rm\chi e}$.
+The $\rm DL$ 90% allowed C.I.
+in typical low and high $\rm m_{\chi}$ ranges $\left[\rm m_{\chi}{=}(100{\pm}10)~{}{\rm and}~{}(1{\pm}0.1)~{}{\rm GeV}\right]$ are depicted in Figures 6b&c, respectively, superimposed with the exclusion contours of the null AM data.
+The complementary roles of the experiments in probing the parameter space can be seen.
+IV Conclusion
+Our ${\Gamma}_{tot}^{2\chi}$ analysis and results indicate that the DL AM allowed regions associated with $\rm\chi e$ are probed and rejected by the various null AM experiments.
 The CDEX [7] data placed more severe constraints on $\sigma^{\rm LR}_{\rm\chi e}$ where the $(1/q^{2})$ dependence favors low threshold experiments.
 The XMASS [6] experiment is more sensitive to $\sigma^{\rm SR}_{\rm\chi e}$ due to its large exposure.
-It can be projected from the results that the DL allowed regions in generic two-WIMP models are ruled out by the null AM experiments.
-Very little room is left to account for the DL AM data with WIMP-induced effects.
+It can be projected that the DL allowed regions in generic two-WIMP models interacting via the $\chi{\rm N}$ and $\rm\chi e$ channels independently are ruled out by the null AM experiments.
+Very little room is left to account for the DL AM data with WIMP-induced $\chi{\rm N}{+}\rm\chi e$ effects.
+V Acknowledgments
 This work is supported by contracts 106-2923-M-001-006-MY5, 110-2112-M-001-029-MY3, and 113-2112-M-001-053-MY3 from the National Science and Technology Council, Taiwan, grants 2021-22/TG2.1 from the National Center of Theoretical Sciences, Taiwan.